# Hallucination of Multimodal Large Language Models: A Survey

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: Preprintccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learning
@@ 1. Introduction
Recently, the emergence of large language models (LLMs) (OpenAI, 2022; Google, 2023; Touvron et al., 2023; Penedo et al., 2023; Zhao et al., 2023c) has dominated a wide range of tasks in natural language processing (NLP), achieving unprecedented progress in language understanding (Hendrycks et al., 2021; Huang et al., 2023a), generation (Zhang et al., 2023b; Zhu et al., 2023b) and reasoning (Wei et al., 2022; Kojima et al., 2022; Qiao et al., 2022; Yu et al., 2023d; Chu et al., 2023). Leveraging the capabilities of robust LLMs, multimodal large language models (MLLMs) (Liu et al., 2023c; Zhu et al., 2023a; Ye et al., 2023a; Dai et al., 2023), sometimes referred to as large vision-language models (LVLMs), are attracting increasing attention. MLLMs show promising ability in multimodal tasks, such as image captioning (Li et al., 2023b), visual question answering (Dai et al., 2023; Liu et al., 2023c), etc. However, there is a concerning trend associated with the rapid advancement in MLLMs. These models exhibit an inclination to generate hallucinations (Li et al., 2023a; Zhou et al., 2023a; Liu et al., 2024), resulting in seemingly plausible yet factually spurious content.
The problem of hallucination originates from LLMs themselves. In the NLP community, the hallucination problem is empirically categorized into two types (Huang et al., 2023c): 1) factuality hallucination emphasizes the discrepancy between generated content and verifiable real-world facts, typically manifesting as factual inconsistency or fabrication; 2) faithfulness hallucination refers to the divergence of generated content from user instructions or the context provided by the input, as well as self-consistency within generated content. In contrast to pure LLMs, research efforts of hallucination in MLLMs mainly focus on the discrepancy between generated text response and provided visual content (Li et al., 2023a; Zhou et al., 2023a; Liu et al., 2024), i.e., cross-modal inconsistency. This difference suggests that studies in LLMs cannot be seemingly transferred to MLLMs. Therefore, there is a growing need to comprehensively survey recent advancements in MLLMs’ hallucination phenomena to inspire new ideas and foster the field’s development.
@@
Comparison with existing surveys. In pursuit of reliable generative AI, hallucination stands out as a major challenge, leading to a series of survey papers on its recent advancements. For pure LLMs, there are several surveys (Zhang et al., 2023c; Huang et al., 2023c), describing the landscape of hallucination in LLMs. In contrast, there are very few surveys on hallucination in the field of MLLMs. To the best of our knowledge, there is only one concurrent work (Liu et al., 2024), a short survey on the hallucination problem of LVLMs. However, our survey distinguishes itself in terms of both taxonomy and scope. We present a layered and granular classification of hallucinations, as shown in Fig. 1, drawing a clearer landscape of this field. Additionally, our approach does not limit itself to specific model architectures as prescribed in the work of (Liu et al., 2024), but rather dissects the causes of hallucinations by tracing back to various affecting factors. We cover a larger range of literature both in terms of paper number and taxonomy structure. Furthermore, our mitigation strategies are intricately linked to the underlying causes, ensuring a cohesive and targeted approach.
@@
from Data (§3.1), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Data Quantity, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Insufficient Data, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. AMBER (Wang et al., 2023c), LLaVA-RLHF (Sun et al., 2023) , cause_leaf, text width=20em ] ] ] [ Data Quality, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Noisy Data, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. HalluciDoctor (Yu et al., 2023b), LLaVA-1.5 (Liu et al., 2023b) , cause_leaf, text width=20em ] ] [ Lack of Diversity, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. LRV-Instruction (Liu et al., 2023d), HalluciDoctor (Yu et al., 2023b) , cause_leaf, text width=20em ] ] [ Detailed descriptions
@@
(open question), color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. Chen et al. (Chen et al., 2023c), EOS (Yue et al., 2024) , cause_leaf, text width=20em ] ] ] [ Statistic Bias, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Frequent Objects, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. POPE (Li et al., 2023a) , cause_leaf, text width=20em ] ] [ Objects Occurrence, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. LURE (Zhou et al., 2023a), VCD (Leng et al., 2023) , cause_leaf, text width=20em ] ] ] ] [ Hallucination
@@
from Model (§3.2), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Vision Model, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Information Loss, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. HallusionBench (Liu et al., 2023a), AMBER (Wang et al., 2023c) , cause_leaf, text width=20em ] ] [ Feature Bias, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. Tong et al. (Tong et al., 2024) , cause_leaf, text width=20em ] ] ] [ Language Model, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Parametric Knowledge, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. VCD (Leng et al., 2023), Volcano (Lee et al., 2023) , cause_leaf, text width=20em ] ] ] [ Cross-modal Interface, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Inferior Alignment, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. HACL (Jiang et al., 2023), Halle-Switch (Zhai et al., 2023) , cause_leaf, text width=20em ] ] ] ] [ Hallucination
@@
from Training (§3.3), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Sequence Supervision, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. MOCHa (Ben-Kish et al., 2023), OPERA (Huang et al., 2023b) , cause_leaf, text width=20em ] ] [ Visual Supervision, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. Chen et al. (Chen et al., 2023c) , cause_leaf, text width=20em ] ] [ Human Feedback, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. RLHF-V (Yu et al., 2023c) , cause_leaf, text width=20em ] ] ] [ Hallucination from
@@
Inference (§3.4), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Lose Visual Attention, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. OPERA (Huang et al., 2023b), HaELM (Wang et al., 2023e) , cause_leaf, text width=20em ] ] ] ] [ Hallucination
@@
Benchmarks(§4), color=lightplum!100, fill=lightplum!100, text=black [ Hallucination Metrics, color=lightplum!100, fill=lightplum!60, text=black [ CHAIR, color=lightplum!100, fill=lightplum!30, text=black [ CHAIR (Rohrbach et al., 2018) , detect_leaf, text width=20em ] ] [ POPE, color=lightplum!100, fill=lightplum!30, text=black [ POPE (Li et al., 2023a) , detect_leaf, text width=20em ] ] [ LLM-based, color=lightplum!100, fill=lightplum!30, text=black [ e.g. GAVIE (Liu et al., 2023d), HaELM (Wang et al., 2023e), HallusionBench (Liu et al., 2023a) , detect_leaf, text width=20em ] ] [ Others, color=lightplum!100, fill=lightplum!30, text=black [ e.g. Faith-Score (Jing et al., 2023), AMBER (Wang et al., 2023c) , detect_leaf, text width=20em ] ] ] [ Hallucination Benchmarks, color=lightplum!100, fill=lightplum!60, text=black [ Discriminative Task, color=lightplum!100, fill=lightplum!30, text=black [ e.g. POPE (Li et al., 2023a), RAH-Bench (Chen et al., 2023c), FGHE (Wang et al., 2023a) , detect_leaf, text width=20em ] ] [ Generative Task, color=lightplum!100, fill=lightplum!30, text=black [ e.g. GAVIE (Liu et al., 2023d), Faith-Score (Jing et al., 2023) , detect_leaf, text width=20em ] ] ] ] [ Hallucination
@@
Hallucinations (§5.1), color=harvestgold!100, fill=harvestgold!60, text=black [ Introducing
Negative Data, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. LRV-Instruction (Liu et al., 2023d),mitigate_leaf, text width=20em ] ] [ Introducing
Counterfactual Data, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. HalluciDoctor (Yu et al., 2023b),mitigate_leaf, text width=20em ] ] [ Mitigating Noises
and Errors, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. ReCaption (Wang et al., 2023a), EOS (Yue et al., 2024),mitigate_leaf, text width=20em ] ] ] [ Mitigating Model-related
Hallucinations (§5.2), color=harvestgold!100, fill=harvestgold!60, text=black [ Scale-up Resolution, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. LLaVA-1.5 (Liu et al., 2023b), InternVL (Chen et al., 2023a), HallE-Switch (Zhai et al., 2023), mitigate_leaf, text width=20em ] ] [ Versatile
Vision Encoders, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. VCoder (Jain et al., 2023), IVE (He et al., 2024), mitigate_leaf, text width=20em ] ] [ Dedicated Module, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. HallE-Switch (Zhai et al., 2023), mitigate_leaf, text width=20em ] ] ] [ Mitigating Training-related
Hallucinations (§5.3), color=harvestgold!100, fill=harvestgold!60, text=black [ Auxiliary Supervision, color=harvestgold!100, fill=harvestgold!40, text=black [ Visual Supervision, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. Chen et al. (Chen et al., 2023c), mitigate_leaf, text width=20em ] ] [ Contrastive Loss, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. HACL (Jiang et al., 2023),mitigate_leaf, text width=20em ] ] [ Others, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. EOS (Yue et al., 2024),mitigate_leaf, text width=20em ] ] ] [ Reinforcement Learning, color=harvestgold!100, fill=harvestgold!40, text=black [ Automatic
Metric-based, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. MOCHa (Ben-Kish et al., 2023),mitigate_leaf, text width=20em ] ] [ RLAIF-based, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. HA-DPO (Zhao et al., 2023b), POVID (Zhou et al., 2024),mitigate_leaf, text width=20em ] ] [ RLHF-based, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. LLaVA-RLHF (Sun et al., 2023), RLHF-V (Yu et al., 2023c),mitigate_leaf, text width=20em ] ] ] ] [ Mitigating Inference-related
@@ 2.2. Multimodal Large Language Models
MLLMs (Liu et al., 2023c; Zhu et al., 2023a; Ye et al., 2023a; Dai et al., 2023) typically refers to a series of models that enable LLMs to perceive and comprehend data from various modalities. Among them, vision+LLM is particularly prominent, owing to the extensive research on vision-language models (VLMs) (Radford et al., 2021; Jia et al., 2021; Yu et al., [n. d.]) prior to LLMs. As a result, MLLMs are sometimes referred to as vision-LLMs (VLLMs) or large vision language models (LVLMs). The goal of MLLMs is to activate the visual capabilities of LLMs, enabling them to ”see” the world via images or videos. Combined with strong reasoning and language generation abilities, MLLMs trigger a series of downstream tasks in multimodal domains, such as image/video captioning and visual question answering. Additionally, MLLMs serve as the foundation for applications in other fields, such as AI assistants, embodied agents, and robotics.
Integrating the two modalities of vision and language involves primarily two types of approaches. The first line of work is built upon off-the-shelf pre-trained uni-modal models. Specifically, these MLLMs usually incorporate a learnable interface between pre-trained visual encoders and LLMs. The interface extracts and integrates information from visual modalities. Such interfaces can be further categorized into 1) learnable query-based and 2) projection layer based. Learnable query-based methods, exemplified by Q-Former (Li et al., 2023b), as used in MiniGPT-4 (Zhu et al., 2023a) and Instruct-BLIP (Dai et al., 2023), utilize a set of learnable query tokens to capture visual signals via cross-attention. Projection layer-based methods, as widely applied in LLaVA (Liu et al., 2023c), Shikra (Chen et al., 2023b), etc., involve training a linear projection layer or a Multi-Layer Perceptron (MLP) module to transform extracted visual features. Both types of interfaces aim to transform pre-trained visual features into the input space of pre-trained LLMs.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: Preprintccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processingccs: Computing methodologies Machine learning
@@ 1. Introduction
Recently, the emergence of large language models (LLMs) (OpenAI, 2022; Google, 2023; Touvron et al., 2023; Penedo et al., 2023; Zhao et al., 2023c) has dominated a wide range of tasks in natural language processing (NLP), achieving unprecedented progress in language understanding (Hendrycks et al., 2021; Huang et al., 2023a), generation (Zhang et al., 2023b; Zhu et al., 2023b) and reasoning (Wei et al., 2022; Kojima et al., 2022; Qiao et al., 2022; Yu et al., 2023d; Chu et al., 2023). Leveraging the capabilities of robust LLMs, multimodal large language models (MLLMs) (Liu et al., 2023c; Zhu et al., 2023a; Ye et al., 2023a; Dai et al., 2023), sometimes referred to as large vision-language models (LVLMs), are attracting increasing attention. MLLMs show promising ability in multimodal tasks, such as image captioning (Li et al., 2023b), visual question answering (Dai et al., 2023; Liu et al., 2023c), etc. However, there is a concerning trend associated with the rapid advancement in MLLMs. These models exhibit an inclination to generate hallucinations (Li et al., 2023a; Zhou et al., 2023a; Liu et al., 2024c), resulting in seemingly plausible yet factually spurious content.
The problem of hallucination originates from LLMs themselves. In the NLP community, the hallucination problem is empirically categorized into two types (Huang et al., 2023c): 1) factuality hallucination emphasizes the discrepancy between generated content and verifiable real-world facts, typically manifesting as factual inconsistency or fabrication; 2) faithfulness hallucination refers to the divergence of generated content from user instructions or the context provided by the input, as well as self-consistency within generated content. In contrast to pure LLMs, research efforts of hallucination in MLLMs mainly focus on the discrepancy between generated text response and provided visual content (Li et al., 2023a; Zhou et al., 2023a; Liu et al., 2024c), i.e., cross-modal inconsistency. This difference suggests that studies in LLMs cannot be seemingly transferred to MLLMs. Therefore, there is a growing need to comprehensively survey recent advancements in MLLMs’ hallucination phenomena to inspire new ideas and foster the field’s development.
@@
Comparison with existing surveys. In pursuit of reliable generative AI, hallucination stands out as a major challenge, leading to a series of survey papers on its recent advancements. For pure LLMs, there are several surveys (Zhang et al., 2023c; Huang et al., 2023c), describing the landscape of hallucination in LLMs. In contrast, there are very few surveys on hallucination in the field of MLLMs. To the best of our knowledge, there is only one concurrent work (Liu et al., 2024c), a short survey on the hallucination problem of LVLMs. However, our survey distinguishes itself in terms of both taxonomy and scope. We present a layered and granular classification of hallucinations, as shown in Fig. 1, drawing a clearer landscape of this field. Additionally, our approach does not limit itself to specific model architectures as prescribed in the work of (Liu et al., 2024c), but rather dissects the causes of hallucinations by tracing back to various affecting factors. We cover a larger range of literature both in terms of paper number and taxonomy structure. Furthermore, our mitigation strategies are intricately linked to the underlying causes, ensuring a cohesive and targeted approach.
@@
from Data (§3.1), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Data Quantity, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Insufficient Data, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. AMBER (Wang et al., 2023b), LLaVA-RLHF (Sun et al., 2023) , cause_leaf, text width=20em ] ] ] [ Data Quality, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Noisy Data, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. HalluciDoctor (Yu et al., 2023b), LLaVA-1.5 (Liu et al., 2023b) , cause_leaf, text width=20em ] ] [ Lack of Diversity, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. LRV-Instruction (Liu et al., 2023d), HalluciDoctor (Yu et al., 2023b) , cause_leaf, text width=20em ] ] [ Detailed descriptions
@@
(open question), color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. Chen et al. (Chen et al., 2023b), EOS (Yue et al., 2024) , cause_leaf, text width=20em ] ] ] [ Statistic Bias, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Frequent Objects, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. POPE (Li et al., 2023a) , cause_leaf, text width=20em ] ] [ Objects Occurrence, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. LURE (Zhou et al., 2023a), VCD (Leng et al., 2023) , cause_leaf, text width=20em ] ] ] ] [ Hallucination
@@
from Model (§3.2), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Vision Model, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Information Loss, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. HallusionBench (Liu et al., 2023a), AMBER (Wang et al., 2023b) , cause_leaf, text width=20em ] ] [ Feature Bias, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. Tong et al. (Tong et al., 2024) , cause_leaf, text width=20em ] ] ] [ Language Model, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Parametric Knowledge, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. VCD (Leng et al., 2023), Volcano (Lee et al., 2023) , cause_leaf, text width=20em ] ] ] [ Cross-modal Interface, color=lighttealblue!100, fill=lighttealblue!40, text=black [ Inferior Alignment, color=lighttealblue!100, fill=lighttealblue!20, text=black [ e.g. HACL (Jiang et al., 2023), Halle-Switch (Zhai et al., 2023) , cause_leaf, text width=20em ] ] ] ] [ Hallucination
@@
from Training (§3.3), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Sequence Supervision, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. MOCHa (Ben-Kish et al., 2023), OPERA (Huang et al., 2023b) , cause_leaf, text width=20em ] ] [ Visual Supervision, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. Chen et al. (Chen et al., 2023b) , cause_leaf, text width=20em ] ] [ Human Feedback, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. RLHF-V (Yu et al., 2023c) , cause_leaf, text width=20em ] ] ] [ Hallucination from
@@
Inference (§3.4), color=lighttealblue!100, fill=lighttealblue!60, text=black [ Visual Attention Deficiency, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. OPERA (Huang et al., 2023b), HaELM (Wang et al., 2023d), M3ID (Favero et al., 2024) , cause_leaf, text width=20em ] ] [ Trap Visual Tokens, color=lighttealblue!100, fill=lighttealblue!40, text=black [ e.g. AvisC (Woo et al., 2024b), VTI (Liu et al., 2025) , cause_leaf, text width=20em ] ] ] ] [ Hallucination
@@
Benchmarks(§4), color=lightplum!100, fill=lightplum!100, text=black [ Hallucination Metrics, color=lightplum!100, fill=lightplum!60, text=black [ CHAIR, color=lightplum!100, fill=lightplum!30, text=black [ CHAIR (Rohrbach et al., 2018) , detect_leaf, text width=20em ] ] [ POPE, color=lightplum!100, fill=lightplum!30, text=black [ POPE (Li et al., 2023a) , detect_leaf, text width=20em ] ] [ LLM-based, color=lightplum!100, fill=lightplum!30, text=black [ e.g. GAVIE (Liu et al., 2023d), HaELM (Wang et al., 2023d), HallusionBench (Liu et al., 2023a) , detect_leaf, text width=20em ] ] [ Others, color=lightplum!100, fill=lightplum!30, text=black [ e.g. Faith-Score (Jing et al., 2023), AMBER (Wang et al., 2023b) , detect_leaf, text width=20em ] ] ] [ Hallucination Benchmarks, color=lightplum!100, fill=lightplum!60, text=black [ Discriminative Task, color=lightplum!100, fill=lightplum!30, text=black [ e.g. POPE (Li et al., 2023a), MME (Fu et al., 2023), MMBench (Liu et al., 2024a) , detect_leaf, text width=20em ] ] [ Generative Task, color=lightplum!100, fill=lightplum!30, text=black [ e.g. MMHal-Bench (Sun et al., 2023), AMBER (Wang et al., 2023b) , detect_leaf, text width=20em ] ] ] ] [ Hallucination
@@
Hallucinations (§5.1), color=harvestgold!100, fill=harvestgold!60, text=black [ Negative Data, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. LRV-Instruction (Liu et al., 2023d),mitigate_leaf, text width=20em ] ] [ Counterfactual Data, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. HalluciDoctor (Yu et al., 2023b),mitigate_leaf, text width=20em ] ] [ Reasoning Data, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. REVERIE (Zhang et al., 2024c),mitigate_leaf, text width=20em ] ] [ Clean Data, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. ReCaption (Wang et al., 2024d), EOS (Yue et al., 2024),mitigate_leaf, text width=20em ] ] ] [ Mitigating Model-related
Hallucinations (§5.2), color=harvestgold!100, fill=harvestgold!60, text=black [ Scale-up Resolution, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. LLaVA-1.5 (Liu et al., 2023b), InternVL (Chen et al., 2024d), HallE-Switch (Zhai et al., 2023), mitigate_leaf, text width=20em ] ] [ Versatile
Vision Encoders, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. VCoder (Jain et al., 2023), IVE (He et al., 2024a), mitigate_leaf, text width=20em ] ] [ Dedicated Module, color=harvestgold!100, fill=harvestgold!40, text=black [ e.g. HallE-Switch (Zhai et al., 2023), PATCH (Shang et al., 2024), mitigate_leaf, text width=20em ] ] ] [ Mitigating Training-related
Hallucinations (§5.3), color=harvestgold!100, fill=harvestgold!60, text=black [ Auxiliary Supervision, color=harvestgold!100, fill=harvestgold!40, text=black [ Visual Supervision, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. Chen et al. (Chen et al., 2023b), mitigate_leaf, text width=20em ] ] [ Contrastive Learning, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. HACL (Jiang et al., 2023), HALVA (Sarkar et al., 2025),mitigate_leaf, text width=20em ] ] [ Others, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. EOS (Yue et al., 2024), CCA (Xing et al., 2024a),mitigate_leaf, text width=20em ] ] ] [ Reinforcement Learning, color=harvestgold!100, fill=harvestgold!40, text=black [ Automatic
Metric-based, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. MOCHa (Ben-Kish et al., 2023),mitigate_leaf, text width=20em ] ] [ RLAIF-based, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. HA-DPO (Zhao et al., 2023b), POVID (Zhou et al., 2024a), FGAIF (Jing and Du, 2024),mitigate_leaf, text width=20em ] ] [ RLHF-based, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. LLaVA-RLHF (Sun et al., 2023), RLHF-V (Yu et al., 2023c),mitigate_leaf, text width=20em ] ] [ Visual Generative Feedback, color=harvestgold!100, fill=harvestgold!20, text=black [ e.g. ESREAL (Kim et al., 2024c), ConVis (Park et al., 2024),mitigate_leaf, text width=20em ] ] ] ] [ Mitigating Inference-related
@@ 2.2. Multimodal Large Language Models
MLLMs (Liu et al., 2023c; Zhu et al., 2023a; Ye et al., 2023a; Dai et al., 2023; Bai et al., 2024b; Zhao et al., 2024b) typically refers to a series of models that enable LLMs to perceive and comprehend data from various modalities. Among them, vision+LLM is particularly prominent, owing to the extensive research on vision-language models (VLMs) (Radford et al., 2021; Jia et al., 2021; Yu et al., [n. d.]) prior to LLMs. As a result, MLLMs are sometimes referred to as vision-LLMs (VLLMs) or large vision language models (LVLMs). The goal of MLLMs is to activate the visual capabilities of LLMs, enabling them to ”see” the world via images or videos. Combined with strong reasoning and language generation abilities, MLLMs trigger a series of downstream tasks in multimodal domains, such as image/video captioning (Wang et al., 2020; Bai et al., 2021; Fan et al., 2023), visual question answering, cross-modal retrieval (Bai et al., 2025b), and so on. Additionally, MLLMs serve as the foundation for applications in other fields, such as AI assistants, embodied agents, and robotics. Recently, there is a trend of developing unified multimodal models (Xie et al., 2024b; Zhou et al., 2024c; Chen et al., 2025c) that integrates both understanding and generation, which are essentially visual-to-text and text-to-visual generation (Rombach et al., 2022; Bai et al., 2024a, 2025a). In this survey, we focus on the multimodal understanding task, i.e., visual-to-text generation.
Integrating the two modalities of vision and language involves primarily two types of approaches. The first line of work is built upon off-the-shelf pre-trained uni-modal models. Specifically, these MLLMs usually incorporate a learnable interface between pre-trained visual encoders and LLMs. The interface extracts and integrates information from visual modalities. Such interfaces can be further categorized into 1) learnable query-based and 2) projection layer based. Learnable query-based methods, exemplified by Q-Former (Li et al., 2023b), as used in MiniGPT-4 (Zhu et al., 2023a) and Instruct-BLIP (Dai et al., 2023), utilize a set of learnable query tokens to capture visual signals via cross-attention. Projection layer-based methods, as widely applied in LLaVA (Liu et al., 2023c), Shikra (Chen et al., 2023a), etc., involve training a linear projection layer or a Multi-Layer Perceptron (MLP) module to transform extracted visual features. Both types of interfaces aim to transform pre-trained visual features into the input space of pre-trained LLMs.